Evaluate the exponential equation for three positive values of x, three negative values of x, and at x=0
y= -3^-x
If I use -3, -2, -1, 0, 1, 2, 3. I'm confused as to the negative x shows that it automatically becomes 1. Is this correct?
-3^-(-3)
y= 9?

Answer by RAY100(1637)

y = -3^(-x),, or y=1/-3^x
.
x,,,,,,,y
.
0,,,,,,-1,,,,,,,,,,,,,,,,y=-3^(-0) = -1,,,,,,,,,,,,remember,,,3^0 =1
-3,,,,,,,-27,,,,,,,,,,,,,y=-3^(--3)= -3^3=-27
-2,,,,,,,+9,,,,,,,,,,,,,,y=-3^(--2)=-3^2=9
-1,,,,,,,-3,,,,,,,,,,,,,,y=-3^(--1)=-3^1=-3
1,,,,,,,-1/3,,,,,,,,,,,,,y=-3^(-1) =-1/3,,,,,,,,,or y=1/-3^1=-1/3
2,,,,,,,+1/9,,,,,,,,,,,,,y=-3^(-2)=+1/9,,,,,,,,,,or y=1/-3^2=1/9
3,,,,,,,-1/27,,,,,,,,,,,,y=-3^(-3)=-1/27,,,,,,,,,or y=1/-3^3=-1/27
.